OpenStudy (anonymous):

graph the equation by plotting 3 points -3y=5x - 7

OpenStudy (saifoo.khan):

y = -5/3x + 7

OpenStudy (anonymous):

u r wrong saifoo -3 will divide 7 also to give 7/3

OpenStudy (anonymous):

-3y = 5x -7 3y = -5x +7 -5x + 7 y = --------- 3

OpenStudy (anonymous):

put x = -1, we get 5 + 7 y = ------- = 12/3 = 4 3 so first point is (-1,4)

OpenStudy (anonymous):

okhay im confused

OpenStudy (anonymous):

why confused ???

OpenStudy (saifoo.khan):

Oh yeah, my bad. :/

OpenStudy (anonymous):

put x = 2, we get -10 + 7 y = --------- = -3/3 = -1 3 so second point is (2,-1)

OpenStudy (anonymous):

because saifoo idk

OpenStudy (saifoo.khan):

harkiat is correct.

OpenStudy (anonymous):

put x = 5, we get -25 + 7 y = --------- = -18/3 = -6 3 so third point is (5,-6)

OpenStudy (anonymous):

okhay thanks you'll get a medal anyway

OpenStudy (saifoo.khan):

:/